THE VIBRATION-ROTATION EMISSION SPECTRUM OF GASEOUS HZnCl

SHANSHAN YU, \underline{ALIREZA SHAYESTEH}, DEJIAN FU, PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, ON, N2L 3G1, Canada.

Gaseous HZnCl has been synthesized for the first time in a high-temperature tube furnace with a D.C. discharge of a flowing mixture of pure HCl and zinc vapor. The vibration-rotation emission spectrum of gaseous HZnCl was measured at high resolution using a Fourier transform spectrometer. The H−Zn stretching mode (ν_{1}) of the H^{64}Zn^{35}Cl, H^{66}Zn^{35}Cl, H^{68}Zn^{35}Cl, and H^{64}Zn^{37}Cl species, as well as the 2ν_{1} − ν_{1} hot band of the most abundant isotopologue (H^{64}Zn^{35}Cl) were observed near 1966 cm^{−1}. A least-squares fit was performed for each of four observed isotopologues and their spectroscopic constants were determined.